# Predicting function of evolutionarily implausible DNA sequences

### v2

@@ 2.4 Evaluation Metrics
We computed the mean base-pair log-likelihood (LL) as the sequence-level LL score. To account for the distinct pretraining objectives of causal language models (CLMs) and masked language models (MLMs), we applied different LL computation strategies, as detailed in Appendix B.1. After obtaining the LL distributions for both the original cassette and its virtual mutants (Nullsettes), we performed a one-sided paired permutation test to assess whether the mutant cassettes exhibit a significant decrease in LL compared to the original sequence (Figure 1B). Additional methodological details are provided in Appendix B.2.
3 Results
We benchmarked these gLMs and their variants across five datasets, as shown in Appendix C. Representative models were selected for further analysis in the following sections.
3.1 Zero-shot Nullsettes prediction performance deteriorates for sequences with low gLM likelihood
The evolutionary plausibility of a sequence, as predicted by a gLM, strongly correlates with the gLM’s performance in predicting mutation effects for the sequence. Figure 2A compares gLM prediction performance of individual functional cassettes and their predicted LL for two models, Evo-2-7B and NucleotideTransformer-2.5B-MS. Here success rate for each cassette is defined as the proportion of Nullsettes (11 in eukaryote, 19 in prokaryote) whose LL score is lower than the original, nonmutant cassette. For both models, we observe strong correlations between the nonmutant LL score and the gLM’s ability to identify LOF mutants. For example, for the E.coli (Low) dataset (Lagator et al., 2022), zero-shot Nullsettes prediction with NT-2.5B-MS is nearly impossible when the nonmutant LL is below -25, and easily accomplished for nonmutants with LL above -22. In fact, Figure 2B shows that all models, despite variations in architecture and training, exhibits a positive correlation between sequence likelihood and mutant prediction performance.
The Evo-2-7B model achieves the best overall performance in zero-shot Nullsettes prediction. Evaluating zero-shot prediction using all sequences within a dataset (See Methods), Figure 2C shows that Evo-2-7B achieves the best overall performance across all five datasets. Specifically for the E.coli and yeast MPRA data using random promoters (left), which are less evolutionarily plausible than more natural sequences shown on the right (Appendix Figure 4).
Figure 3: Optimal likelihood range for Nullsettes prediction varies with sequence length. A) Length in base pairs and normalized gLM log likelihood of all nonmutant cassettes, averaged across all 12 models. B) Contour of scatter plots similar to those in Figure 2A, colored by dataset, dotted vertical lines are rough estimates of optimal LL value for sequences from different datasets. Human data for Evo 1.5 is cut off for better visual due to large LL score.
3.2 Optimal likelihood range predictive of strong model performance varies with sequence length
Although mutant prediction performance correlates with sequence likelihood, the range of LL scores for which a model performs well depends highly on the length of the DNA sequence. The range of LL scores for which models achieve strong performance is higher for longer sequences, partly due to the fact that gLMs tend to assign higher LL scores to longer cassettes (Figure 3A). For example, Figure 3B shows that the NT-2.5B-MS model achieves strong performance for cassettes with log likelihood above $-22$ for both E. coli datasets, as indicated by the gray dotted line. However, for sequences from the Human GPRA dataset which are $2.5$-folds longer, only LL score above $-17$ achieved strong performance. Similarly sequences from the yeast (low) dataset are $30\%$ longer compared to sequences from the E. coli datasets. Thus, for GPN-promoter and Evo-1.5, the optimal likelihood range is also significant higher for the yeast (low) dataset.
Altogether, there does not appear to be a single optimal likelihood range within which a model would be able to make zero-shot mutation effect predictions, as sequence properties such as length can have a significant effect on optimal likelihood range.
4 Discussion
In this study, we introduce Nullsettes, the first systematic benchmark for assessing the ability of genomic language models (gLMs) to predict loss-of-function mutations in synthetic expression cassettes. By enabling simple in silico translocation of sequences, Nullsettes facilitates generalizable, cassette-independent evaluations of gLM functional performance. Beyond general performance benchmarking, our results reveal a strong correlation between gLM performance on mutation prediction tasks and the model’s predicted likelihood of the original, non-mutant sequence. Moreover, we demonstrate that the optimal likelihood range for accurate mutant discrimination is not fixed but varies significantly with sequence length. Together, we hope this benchmark and its findings will inspire the development of gLMs that move beyond evolutionary priors to support the rational design of functional, out-of-distribution genetic sequences.
Software and Data
The source code for benchmarking gLM on Nullsettes is publicly available at: https://github.com/cellethology/GLM-Nullsette-Benchmark.
Acknowledgements
The authors thank members of the Cell Ethology Lab for valuable comments and suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of genomic language model and its application in synthetic biology. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Datasets
A.1 Prokaryote
To evaluate the regulatory grammar of prokaryotes on genomic language models, we utilized two benchmark datasets derived from synthetic expression constructs in Escherichia coli. The first, the Kosuri dataset (Kosuri et al., 2013), comprises combinatorial assemblies of promoters and ribosome binding sites (RBS) upstream of a superfolder GFP (sfGFP) reporter, designed to dissect transcriptional and translational contributions to gene expression. The second, the Lagator dataset (Lagator et al., 2022), features a large-scale library of random promoter sequences with experimentally measured activity, enabling evaluation of models’ ability to generalize to highly diverse, out-of-distribution regulatory inputs.
A.2 Eukaryote
To evaluate the capacity of genomic language models to generalize to eukaryotic regulatory logic, we utilized two large-scale massively parallel reporter assay (MPRA) datasets. The Zahm dataset (Zahm et al., 2024) comprises a library of 6,144 synthetic promoters constructed by combining transcriptional response elements (TREs) from 229 human and mouse transcription factors with minimal promoters. These constructs were assayed across multiple human cell lines and stimulatory conditions to quantify dynamic, stimulus-specific gene expression. The deBoer dataset (de Boer et al., 2020) includes a comprehensive collection of 100 million randomized promoter sequences in yeast, enabling high-resolution dissection of cis-regulatory grammar under out-of-distribution scenario.
A.3 Promoter selection
A.3.1 Kosuri and Lagator datasets
For the Kosuri dataset (Kosuri et al., 2013), we selected candidate promoter–RBS pairs with protein expression levels exceeding $\mu+1.5\sigma$ (10,165) across the full distribution. In contrast, given the limited number of highly active constructs in the Lagator dataset, we directly selected the top 1,500 promoter sequences ranked by protein output. We then compared the log-likelihood (LL) distributions of promoter–RBS groups using genomic language models including Evo1 8k, GENERator 3B, GPN, and Nucleotide Transformer 2.5B multispecies. As the Lagator dataset (Lagator et al., 2022) consists of randomly generated promoters, whereas the Kosuri library is more rationally designed, we prioritized Kosuri promoter–RBS pairs whose LL distributions surpassed those from Lagator (Appendix Figure 4(a)), yielding 1,500 expression cassettes from each dataset.
A.3.2 deBoer dataset
deBoer dataset (de Boer et al., 2020) contains two promoter libraries: Abf1TATA and pTpA. Abf1TATA is designed by embedding conserved transcription factor binding sites such as Abf1 and a canonical TATA box, mimicking features of natural yeast promoters; while pTpA consists of synthetic promoters constructed with a simple poly-T–poly-A architecture, lacking specific transcription factor motifs and serving as a minimal, randomized control library. Similar to Kosuri vs Lagator, we selectively constructed 1,500 active promoters for each dataset based on LL distribution where pTpA is left-shifted comparing to Abf1TATA as Abf1TATA is more conservative (Appendix Figure 4(b)). The source can be found at NCBI’s GEO: GSE104878111https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE104878.
A.3.3 Zahm dataset
The Zahm dataset (Zahm et al., 2024) comprises synthetic promoters built by coupling one of three minimal promoters—minCMV, minProm, or minTK—with diverse transcriptional response element (TRE) units. To construct a representative set of mammalian expression cassettes, we selected the top 1,500 promoter-TRE combinations exhibiting the highest transcriptional activity. The source can be found at NCBI’s GEO: GSE271608222https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE271608.
(a) Kosuri vs Lagator
(b) deBoer TATA vs deBoer pTpA
Figure 4: Comparison of promoter log-likelihood distribution
A.4 Virtual mutant
Translocation mutants were systematically engineered by permuting the canonical order of key regulatory and coding elements—promoter, ribosome binding site (RBS), start codon, coding sequence (CDS), stop codon, and terminator in prokaryotic systems; and promoter, start codon, CDS, stop codon, and terminator in eukaryotic systems. Each permutation aims to disrupt the natural transcription-translation flow, creating syntactically or functionally invalid constructs. A total of 19 translocation variants were designed for prokaryotic systems and 11 for eukaryotic systems. All constructs were rendered completely non-functional by design. All translocation combinations can be found in Appendix Table 3 for prokaryote and Appendix Table 4 for eukaryote.
Appendix B Evaluation Metrics
B.1 Mean Log-likelihood
@@
For casual language models (CLM), the log-likelihood of a nucleic acid sequence $X=(x_{1},x_{2},...,x_{n})$ is computed using logits output from the model. Since CLM operates in an autogressive manner, each token $x_{t}$ is predicted based on all previous token $x<t$. Given the model’s ${logits}_{t}$ at each position $t$, the probability of the ground-truth token is obtained via softmax:
@@
Masked language models (MLMs) differ from autoregressive models in their inference behavior, as they are not inherently designed for sequential generation. To address this, Wang & Cho (2019) introduced the pseudo log-likelihood (PLL) approach, wherein each token in a sequence is masked individually to compute token-wise conditional probabilities. More recently, Gordon et al. (2024) proposed Single-Inference Pseudo Log-Likelihood, an efficient approximation that enables linear-time inference for BERT-style MLMs by exploiting training-time masking dynamics.
@@ B.2 One-sided Paired Permutation Test
Appendix C Model Performance Evaluation
While representative models were selected for the main analysis, we conducted an extensive evaluation of various model versions across all datasets. As shown in Appendix Figure 5, the Evo series models (Nguyen et al., 2024; Merchant et al., 2024; Brixi et al., 2025) consistently outperformed all others in the Nullsettes tasks, with all the Evo variants ranking at the top. This advantage likely reflects the superior architectural design and scaling strategy employed by the Evo models. Among non-Evo models, METAGENE-1 (Liu et al., 2025b) achieved the highest performance, likely due to its training on over 1.5 trillion base pairs of DNA and RNA sequences derived from wastewater metagenomes. The unique breadth and diversity of metagenomic data may provide this model with an enhanced ability to generalize across sequence contexts. The Nucleotide Transformer (NT) (Dalla-Torre et al., 2024) series followed, ranking third overall. These findings highlight the impact of both model architecture and training data diversity on generalization performance across regulatory sequence prediction tasks.
@@ Appendix C Model Performance Evaluation
Figure 5: The number of cumulative successful predictions made by each model series across four datasets: Abf1TATA and pTpA (from deBoer), Zahm, Kosuri, and Lagator.
Nucleotide Transformer 2.5B multispecies (Dalla-Torre et al., 2024) 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e. roughly 29B tokens. MLM 6-mers Transformer, 2.5B params 1000 bp
Nucleotide Transformer 500M multispecies (Dalla-Torre et al., 2024) 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e. roughly 29B tokens. MLM 6-mers Transformer, 500M params 1000 bp
Nucleotide Transformer 500M Human ref (Dalla-Torre et al., 2024) GRCh38 human reference genome, resulting in 3B nucleotides, i.e. roughly 500M 6-mer tokens. MLM 6-mers Transformer, 500M params 1000 bp
Evo1 8k (Nguyen et al., 2024) OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context
Evo1 131k (Nguyen et al., 2024) OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 131,072 context using Evo1 8k as the base model
Evo1.5 (Merchant et al., 2024) 50% increase in training data comparing to Evo1 CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context using Evo1 8k as the base model
Evo2 1b base (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 1B params Pretrained with 8192 context length
Evo2 7b base (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 8192 context length
Evo2 7b (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 1M context using Evo2 7b base as the base model
GENERator 3b (Wu et al., 2025) Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate. CLM 6-mer Llama-based decoder, 3B params Pretrained with a context length of 98k bp
GENERator 1.2b (Wu et al., 2025) Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate. CLM 6-mer Llama-based decoder, 1.2B params Pretrained with a context length of 98k bp
METAGENE-1 (Liu et al., 2025b) 1.5T base pairs of DNA and RNA sequences from human wastewater samples. CLM Byte-pair encoding Llama-2, 7B params Pretrained with 512 sequence length
Table 1: Genomic language model details
Caduceus-ph (Schiff et al., 2024) Human reference genome (35M tokens / nucleotide base pairs). MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (reverse complement augmentation)
Caduceus-ps (Schiff et al., 2024) Human reference genome (35M tokens / nucleotide base pairs). MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (No reverse complement augmentation)
GENA-LM human (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly. MLM Byte-pair encoding BERT/BigBird, 110M params About 4,500 nucleotides (512 BPE tokens)
GENA-LM multi (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly augmented by sampling mutations from 1000-genome SNPs (gnomAD dataset) + Multispecies genomes from ENSEMBL release 108. MLM Byte-pair encoding BERT/BigBird, 110M params About 4,500 nucleotides (512 BPE tokens)
GENA-LM large (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly. MLM Byte-pair encoding BERT/BigBird, 336M params About 4,500 nucleotides (512 BPE tokens)
GPN (Benegas et al., 2023) Brassicales reference genome from NCBI Genome. Took the union of exons (with a small intronic flank), promoters (1,000 bp upstream of transcription start sites) as well as an equivalent amount of random windows from the whole genome. MLM Single-nucleotide tokens CNN 512 bp
GPN-Promoter (Benegas et al., 2025a) Animal promoter, genomes of 434 animal species. MLM Single-nucleotide tokens ByteNet, 152M Pretrained on 512 bp sequences centered at TSSs of protein-coding genes
DNABERT 2 (Zhou et al., 2023) Human genome dataset (2.75B nucleotide bases) + Multispecies genome dataset (from 135 species, spread across 6 cetogories). The dataset includes 32.49B nucleotides bases, excluding all sequence with N and retain only ATCG. MLM Byte-pair encoding BERT, 117M params Pretrained on 700 bp length sequences
PDLLM-DNAMamba-6mer (Liu et al., 2025a) Plant reference genomes CLM 6-mer SSM, 130M params Pretrained max token length is 512
Table 2: Genomic language model details (continued)
@@
Table 3: Prokaryotic virtual mutant cases. Translocation-3,4,5 cannot be compensated by a later start codon as there are no other in-frame start codon.
@@
Table 4: Eukaryotic virtual mutant cases. Translocation-2,3,4 cannot be compensated by a later start codon as there are no other in-frame start codon.

### v3

@@ 4.4 Evaluation Metrics
We computed the mean base-pair log-likelihood (LL) as the sequence-level LL score. To account for the distinct pretraining objectives of causal language models and masked language models, we applied different LL computation strategies.
@@
Specifically, for casual language models (CLM), the log-likelihood of a nucleic acid sequence $X=(x_{1},x_{2},...,x_{n})$ is computed using logits output from the model. Since CLM operates in an autogressive manner, each token $x_{t}$ is predicted based on all previous token $x<t$. Given the model’s ${logits}_{t}$ at each position $t$, the probability of the ground-truth token is obtained via softmax:
@@
Masked language models (MLMs) differ from autoregressive models in their inference behavior, as they are not inherently designed for sequential generation. To address this, Wang and Cho [24] introduced the pseudo log-likelihood (PLL) approach, wherein each token in a sequence is masked individually to compute token-wise conditional probabilities. More recently, Gordon et al. [25] proposed Single-Inference Pseudo Log-Likelihood, an efficient approximation that enables linear-time inference for BERT-style MLMs by exploiting training-time masking dynamics.
@@
5 Data availability
The expression cassette of Kosuri dataset [16] was obtained from https://www.addgene.org/47441/. The Lagator dataset [14] was downloaded from https://github.com/szarma/Thermoters. The deBoer dataset [15] was downloaded from https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE104878. The Zahm dataset [17] was downloaded from https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE271608.
6 Code availability
The source code for benchmarking gLMs and the Nullsettes used in this work is publicly available at: https://github.com/cellethology/GLM-Nullsette-Benchmark.
7 Acknowledgements
The authors thank members of the Cell Ethology Lab for valuable comments and suggestions. This work was supported by the Westlake Fellows Program at Westlake University.
8 Author Contributions
Z.J.W. conceived the initial proposal. S.J. and Z.J.W. jointly developed the overall concept and contributed to the framework design. S.J. implemented the framework. S.J., X.L., and Z.J.W. conducted and analyzed the computational experiments. S.J. and Z.J.W. wrote the initial draft of the manuscript.
9 Competing interests
The authors declare no competing interests.
Supplementary Information
Supplemental Figures
Figure S1: Comparison of promoter log-likelihood distribution
@@ 9 Competing interests
Figure S2: The number of cumulative successful predictions made by each model series across four datasets: Abf1TATA and pTpA (from deBoer), Zahm, Kosuri, and Lagator.
Figure S3: Number of Nullsette mutations successfully identified for gLMs with small and large variants, including Evo2 (1B vs. 7B), GENERator (1.2B vs. 3B), NT (500M vs. 2.5B, multispecies), and gLM2 (150M vs. 650M).
Supplemental Tables
@@
Table S1: Prokaryotic virtual mutant cases. Translocation-3,4,5 cannot be compensated by a later start codon as there are no other in-frame start codon.
@@
Table S2: Eukaryotic virtual mutant cases. Translocation-2,3,4 cannot be compensated by a later start codon as there are no other in-frame start codon.
Model series Representative variant
METAGENE1 [26] METAGENE-1 (METAGENE1)
Nucleotide Transformer [27] nucleotide-transformer-2.5b-multi-species (NT-2.5B-NT)
GENERator [28] GENERator-eukaryote-3b-base (GENERator-3B)
GENERanno [18] GENERanno-prokaryote-0.5b-base (GENERanno-0.5B)
Evo1 [29, 30] evo-1.5-8k-base (Evo1.5)
Evo2 [31] evo2_7b_base (Evo2-7B)
DNABERT2 [32] DNABERT-2-117M (DNABERT2)
Caduceus [33] caduceus-ph_seqlen-131k_d_model-256_n_layer-16 (Caduceus)
GPN [8] gpn-brassicales (GPN)
GPN-Promoter [34] gpn-animal-promoter (GPN-Promoter)
gLM2 [35] gLM2_650M (gLM2)
PDLLM [36] PDLLM-dnamamba-6mer (PDLLM)
Table S3: Representative model variant. For each model series, we list a representative variant used in major benchmarking. The official model identifier is listed, followed in parentheses by the shorthand name used throughout the main text. For example, the variant “nucleotide-transformer-2.5b-multi-species” is referred to as “NT-2.5B-NT” in the text.
Evo1 8k [29] OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context
Evo1 131k [29] OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 131,072 context using Evo1 8k as the base model
Evo1.5 [30] 50% increase in training data comparing to Evo1 CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context using Evo1 8k as the base model
Evo2 1B base [31] OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 1B params Pretrained with 8192 context length
Evo2 7B base [31] OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 8192 context length
Evo2 7B [31] OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 1M context using Evo2 7b base as the base model
GENERator eukaryote 3B [28] Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate. CLM 6-mer Llama-based decoder, 3B params Pretrained with a context length of 98k bp
GENERator eukaryote 1.2B [28] Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate. CLM 6-mer Llama-based decoder, 1.2B params Pretrained with a context length of 98k bp
METAGENE-1 [26] 1.5T base pairs of DNA and RNA sequences from human wastewater samples. CLM Byte-pair encoding Llama-2, 7B params Pretrained with 512 sequence length
PDLLM-DNAMamba [36] Plant reference genomes CLM 6-mer/5-mer/4-mer/3-mer/2-mer/single/BPE SSM, 130M params Pretrained max token length is 512
Table S4: Casual language modeling (CLM) based genomic language models
Nucleotide Transformer 2.5B multispecies [27] 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e. roughly 29B tokens. MLM 6-mers Transformer, 2.5B params 1000 bp
Nucleotide Transformer 500M multispecies [27] 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e. roughly 29B tokens. MLM 6-mers Transformer, 500M params 1000 bp
Nucleotide Transformer 500M Human ref [27] GRCh38 human reference genome, resulting in 3B nucleotides, i.e. roughly 500M 6-mer tokens. MLM 6-mers Transformer, 500M params 1000 bp
GENERanno prokaryote 0.5B base [18] 715 billion base pairs of prokaryotic DNA MLM single-nucleotide Transformer encoder, 500M params 8k bp
GENERanno eukaryote 0.5B base [18] 386 billion base pairs of eukaryotic DNA MLM single-nucleotide Transformer encoder, 500M params 8k bp
Caduceus-ph [33] Human reference genome (35M tokens / nucleotide base pairs). MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (reverse complement augmentation)
Caduceus-ps [33] Human reference genome (35M tokens / nucleotide base pairs). MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (No reverse complement augmentation)
GPN [8] Brassicales reference genome from NCBI Genome. Took the union of exons (with a small intronic flank), promoters (1,000 bp upstream of transcription start sites) as well as an equivalent amount of random windows from the whole genome. MLM Single-nucleotide tokens CNN 512 bp
GPN-Promoter [34] Animal promoter, genomes of 434 animal species. MLM Single-nucleotide tokens ByteNet, 152M Pretrained on 512 bp sequences centered at TSSs of protein-coding genes
DNABERT2 [32] Human genome dataset (2.75B nucleotide bases) + Multispecies genome dataset (from 135 species, spread across 6 cetogories). The dataset includes 32.49B nucleotides bases, excluding all sequence with N and retain only ATCG. MLM Byte-pair encoding BERT, 117M params Pretrained on 700 bp length sequences
gLM2 [35] OMG: encodes genomic scaffolds with both amino-acid (CDS) and DNA tokens (315B tokens). MLM Char-level tokens (DNA: lowercase, AA: uppercase) Transformer, 650M params / 150M params Pretrained with a 4096 token context window
Table S5: Mask language modeling (MLM) based genomic language models